OpenStudy (anonymous):

Find the surface area of a cone with a radius of 2.7 cm and a slant height of 8.3 cm, to the nearest square unit. Use 3.14 for p. a. 116 cm2 b. 164 cm2 c. 93 cm2 d. 58 cm2

OpenStudy (some.random.cool.kid):

Alright so what our area?

OpenStudy (anonymous):

@some.random.cool.kid you think you can help me with this problem

OpenStudy (some.random.cool.kid):

I can certainly try! :D

OpenStudy (anonymous):

alright @some.random.cool.kid

OpenStudy (some.random.cool.kid):

A=πr(r+h2+r2)

OpenStudy (some.random.cool.kid):

you have your height as 8.3

OpenStudy (some.random.cool.kid):

so fill this in A=πr(r+h2+r2)=π·2.7·(2.7+8.32+2.72)

OpenStudy (anonymous):

ok

OpenStudy (some.random.cool.kid):

well first you are is A≈96.94cm²

OpenStudy (anonymous):

oh yeah ok

OpenStudy (anonymous):

oh yeah ok

OpenStudy (anonymous):

ok

OpenStudy (some.random.cool.kid):

alright so do you think you can try the next step on your own? I will be trying to solve it while you do k.

OpenStudy (anonymous):

alright

OpenStudy (some.random.cool.kid):

explanation r = 2.7 m h = 8.3 m s = 8.72812 m V = 63.3628 m3 L = 74.0345 m2 B = 22.9022 m2 A = 96.9367 m2 I believe you multiply this to get part however 96.94 x 8.3 wont give you your answer their is another step I believe. radius x height = 22.41 I belive your closet answer is 93cm^2 But can also be 164cm^2 so please challenge me if im wrong but its between both of those answers.

OpenStudy (some.random.cool.kid):

@Nnesha or @nincompoop you guys might able to explain this better than I ...

OpenStudy (some.random.cool.kid):

got it i believe your answer is 93

OpenStudy (some.random.cool.kid):

I used woldfram and found the answer to the area I was confused at first but then figured it out.

OpenStudy (some.random.cool.kid):

hope this helped @michael305